OpenStudy (anonymous):

there are 6 girls and 4 boys in a class. what is the probability that a student chosen at random is a 12 year old girl

OpenStudy (anonymous):

60% because in total there are 10 people

thomaster (thomaster):

You cannot tell because you don't know the ages of the 10 people xD You can say it's a 60% chance that it will be a girl but not age 12

OpenStudy (anonymous):

a 0.18 b 0.3 c 0.45 d 0.9

OpenStudy (ash2326):

We should account that a girl has to be chosen first, then only she can be 12 or not Probability of choosing a girl \[P(G)=\frac{6}{10}=0.6\] probability that she is 12 years old \[P(12)=0.5\] Probability of choosing a 12 year girl \[P=P(G)\times P(12)=0.6\times 0.5=0.3\]

OpenStudy (anonymous):

@ash2326 yes :)

OpenStudy (ash2326):

@edensleka do you get this?

OpenStudy (ash2326):

@perurabo yes, you can. Start it as a new post.

OpenStudy (anonymous):

p=\frac{ 6 }{ 10 }=\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

p=3/5

OpenStudy (kropot72):

The distribution of ages among the girls is not given in the question. If there were no 12 year old girls, the probability of choosing a 12 year old girl would be zero. If all the girls were 12 years old, the probability of choosing a 12 year old girl would be 0.6. Therefore there is not enough information in the question to enable an answer to be given.